If the child’s acceleration is 0.31m/s^2, what is the parent’s acceleration?


1
Expert's answer
2021-02-04T11:52:19-0500

If "F1 = F2," then

"m1*a1 = m2*a2,"

where m - the mass, a - the acceleration.

"a2 = m1*a1\/m2"

So, we must know child's and parent's masses to calculate parent's acceleration.